# Testing LRD in the spectral domain for functional time series in manifolds

### v3

@@ Abstract
keywords:
and
Contents
- 1 Introduction
- 2 Preliminary results under SRD
- 3 Second and fourth order bias asymptotics under LRD
- 4 A test for LRD in functional time series on
- 5 Simulation study
- 6 Final comments and discussion
- A Proofs of the results in Section 2
- B Proofs of the results in Section 3
- C Proofs of the results in Section 4
@@ 1 Introduction
Spherical functional time series analysis helps in understanding the dynamics and spatiotemporal patterns of data that are embedded into the sphere, providing valuable insights for prediction, monitoring, and decision-making. Time series analysis of global temperature data distribution among other climate variables, usually arising in Climate Science and Meteorology, can be performed in a more efficient way by adopting a functional time series framework (see Shah et al. (2024)). That is the case of ocean currents, and other marine functional time series to be analyzed in Oceanography studies (see, e.g., Tirelli et al. (2024); Yarger, Stoev and Hsing (2022)). Other areas demanding this type of techniques are Geophysics, Astronomy and Astrophysics. In the last few decades, the cosmic microwave background radiation variation analysis over time has gained special attention (see Duke et al. (2024); Marinucci (2004); Marinucci and Peccati (2011)). In a more general manifold setting, functional time series analysis is often applied in Medical Imaging, Computer Vision and Graphics (see Wu, Huang and Srivastava (2024); Zhou, Li and Zhu (2013); Zhu et al. (2009), among others). This paper focuses on the spectral analysis of functional time series in manifolds, with special attention to LRD analysis.
@@
The spectral analysis of functional time series has mainly been developed under Short Range Dependence (SRD). In this context, based on the weighted periodogram operator, a nonparametric framework is adopted in Panaretos and Tavakoli (2013a). Particularly, the asymptotic normality of the functional discrete Fourier transform (fDFT), and the weighted periodogram operator of the curve data are proved under suitable summability conditions on the norm of the cumulant spectral density operators. The consistency of the weighted periodogram operator, in the Hilbert–Schmidt operator norm, is derived under SRD. In Panaretos and Tavakoli (2013b), a harmonic principal component analysis of functional time series in the temporal functional spectral domain is also obtained, based on a Karhunen–Loéve–like decomposition, the so–called Cramér–Karhunen–Loéve representation (see also Rubín and Panaretos (2020b)). In the context of functional regression, some applications are presented in Pham and Panaretos (2018), Rubín and Panaretos (2020a) and Tavakoli (2014). Hypothesis testing for detecting modelling differences in functional time series dynamics is addressed in Tavakoli and Panaretos (2016) in the functional spectral domain.
@@
In LRD analysis of functional time series several problems still remain open. One of the key approaches in the current literature is presented in Li, Robinson and Shang (2019), where the eigendecomposition of the long-run covariance operator is considered, under an asymptotic semiparametric functional principal component framework. The consistent estimation of the dimension and the orthonormal functions spanning the dominant subspace, where the projected curve process displays the largest dependence range is derived. Fractionally integrated functional autoregressive moving averages processes constitute an interesting example of this modelling framework.
@@
A first attempt to characterize LRD in functional time series in the spectral domain can be found in Ruiz–Medina (2022), adopting the theoretical framework of operator–valued random fields, including fractional Brownian motion with operator–valued Hurst coefficient (see, e.g., Düker (2018); Düker (2020); Marinucci, Rossi and Vidotto (2020); Rackauskas and Suquet (2010) and Rackauskas and Suquet (2011)). The eigenvalues of the LRD operator are parameterized. These eigenvalues induce different levels of singularity at zero frequency, corresponding to different levels of temporal persistent of the process projected into different eigenspaces of the Laplace Beltrami operator. Under this LRD scenario, the integrated periodogram operator is proved to be asymptotically unbiased in the Hilbert–Schmidt operator norm. Minimum contrast estimation of the LRD operator is achieved in the spectral domain in a weak–consistent way under a Gaussian scenario. Interesting examples of this setting are analyzed in Ovalle–Muñoz and Ruiz–Medina (2024a), where the spectral analysis of multifractionally integrated functional time series in manifolds is considered. In particular, multifractionally integrated spherical functional ARMA models (i.e., multifractionally integrated SPHARMA models) are analyzed through simulations. In this modelling framework, SRD and LRD can coexist at different spherical scales.
@@
Up to our knowledge, no further developments have been achieved in the spectral analysis of LRD functional time series. In this paper we perform a weighted periodogram operator based analysis, requiring the asymptotic analysis of the bias of the integrated empirical fourth–order cumulant spectral density operators, to prove consistency of the integrated weighted periodogram operator under LRD. Its application to spectral statistical hypothesis testing of LRD in –valued correlated sequences constitutes one of the main goals of this work. Here, denotes the space of real–valued square integrable functions on a Riemannian manifold embedded into given by a connected and compact two–point homogeneous space. The topological dimension of is and denotes the normalized Riemannian measure on In what follows, we will consider to be a functional sequence such that for every with denoting the probability measure defined on the basic probability space i.e., for every
@@
The invariance of the elements of the spectral density operator family of under the group of isometries of the manifold is assumed along the paper. A frequency–varying eigenvalue sequence then characterizes the pure point spectra of the elements of the spectral density operator family, with respect to the orthonormal basis of eigenfunctions of the Laplace–Beltrami operator. This invariance assumption is exploited in the derived Central Limit Theorem that characterizes the asymptotic distribution of the proposed test statistic operator under the null hypothesis, which states that displays SRD. In our formulation of the alternative hypothesis on LRD, we adopt a semiparametric framework in terms of a functional parameter given by the LRD operator. In contrast with the approach presented in Ruiz–Medina (2022), here we do not assume a parameterization of the eigenvalues of the LRD operator. Under this scenario, the rate of convergence to zero, in the corresponding norm, of the integrated empirical second and fourth order cumulant spectral density operators is respectively obtained in Lemmas 3.1 and 3.3 under suitable conditions. Proposition 4.1 shows the divergence, in the Hilbert–Schmidt operator norm, of the mean of the test statistic operator under the alternative hypothesis. Theorem 4.2 derives suitable conditions, in particular, on the nonparametric functional spectral factor, to ensure consistency of the integrated weighted periodogram operator in the Hilbert–Schmidt operator norm under LRD. Theorem 4.5 then provides the almost surely divergence of the test statistic in the Hilbert–Schmidt operator norm under the alternative, yielding the consistency of the test.
@@
Theorem 4.5 also plays a crucial role in the implementation in practice of the proposed testing procedure, based on rejecting the null hypothesis when the random Fourier coefficients of the test operator statistic, suitably standardized according to Theorem 2.2, cross an upper or lower tail standard normal critical value. The orthonormal basis involved in the computation of these coefficients is constructed by the tensor product of the eigenfunctions of the Laplace Beltrami operator. The random projection methodology (see Theorem 4.1 in Cuesta-Albertos, Fraiman and Ransford (2007)) can be implemented here to alleviate the dimensionality problem. Specifically, when the moments of our test statistic operator satisfy the Carleman condition under the null hypothesis, our testing procedure is equivalent to rejecting the null hypothesis when the absolute value of a random projection of the test statistic is larger than an upper tail standard normal critical value. In the implementation of the random projection methodology, the Karhunen–Loéve expansion in Lemma 4.6 below can be considered for generation of the involved Gaussian random directions. In the simulation study undertaken, robust empirical sizes, and competitive values of the empirical power are displayed by our testing approach (see Section 5.4).
@@
The outline of the paper is as follows. In Section 1.1, the functional spectral background material is introduced. Our hypothesis testing procedure is formulated in a functional semiparametric spectral framework in Section 1.2. The asymptotic Gaussian distribution of the test statistic operator under the null hypothesis is obtained in Theorem 2.2 in Section 2. Asymptotics of the bias of the integrated empirical second and fourth order cumulant spectral density operators under LRD are derived in Section 3. Section 4 provides the preliminary results required for consistency of the test, which is derived in Theorem 4.5 of this section. Practical implementation is also discussed in Section 4. In Section 5.1, a simulation study is undertaken to illustrate the asymptotic Gaussian distribution of the proposed test statistic operator under the null hypothesis, in the context of SRD spherical functional time series. The consistency of the test is also illustrated in Section 5.2, in the framework of multifractionally integrated spherical functional time series. This numerical analysis is extended in Section 5.3 to a wider family of LRD operators allowing stronger persistency in time, displayed by the projected process in the dominant subspace. Section 5.4 analyzes empirical size and power properties of the test. Section 6 summarizes conclusions of the simulation study, focusing on the large functional sample size properties of our test statistic operator under different bandwidth parameter scenarios, from additional numerical results. The proofs of the results of this paper can be found in the Appendix.
@@ 1.1 Background
Along this work we will assume that in (1) is a stationary zero–mean functional sequence, with nuclear covariance operator family satisfying for every The elements of this family are characterized in the spectral domain by the spectral density operator family The assumed invariance of the elements of these families with respect to the group of isometries of yields to their diagonal series expansion in terms of with being the orthonormal basis of eigenfunctions of the Laplace–Beltrami operator on (see, e.g., Giné (1975); Helgason (1959)). In particular,
where, for every denotes the dimension of the eigenspace associated with the eigenvalue of the Laplace Beltrami operator (see, e.g., Section 2.1 in Ma and Malyarenko (2020)). The equality means identity in the norm of the space of Hilbert–Schmidt operators on the space of complex–valued square integrable functions on Specifically, the equality in (2) means that
@@ 1.2 Hypothesis testing
Stationary SRD functional time series are characterized by the summability of the series of trace norms of the elements of the family of covariance operators (see, e.g., Panaretos and Tavakoli (2013a)). That is, displays SRD if and only if where denotes the space of trace operators on In our setting, this condition can be formulated as follows:
When (8) fails, is said to display LRD. In what follows, we will adopt the LRD scenario introduced in Ruiz–Medina (2022), given by
where the invariant positive self–adjoint operators and are composed yielding the definition of Specifically, denotes the LRD operator on Operator in (9) is interpreted as in Characiejus and Räckauskas (2014), Rackauskas and Suquet (2010) and Rackauskas and Suquet (2011), where plays the role of operator–valued Hurst coefficient in the setting of fractional Brownian motion introduced in this framework. Moreover, is the regular spectral operator reflecting markovianess when the null space of coincides with Specifically, satisfies
where the operator integrals are understood as improper operator Stieltjes integrals which converge strongly (see, e.g., Section 8.2.1 in Ramm (2005)).
We will apply the spectral theory of self–adjoint operators (see, e.g., Dautray and Lions (1985)) in terms of the common spectral kernel
@@
where denotes the sequence of positive eigenvalues. For each is a continuous positive slowly varying function at in the Zygmund’s sense (see Definition 6.6 in Beran (2017), and Assumption IV in Ruiz–Medina (2022)). Equation (10) can be equivalently expressed, in terms of as
@@
for every (see Gelfand and Vilenkin (1964) for the usual notion of convergence in the sense of generalized functions or distributions). Here, denotes the Dirac Delta distribution at zero frequency. Hence, in what follows we adopt the notation representing a truncated Dirac Delta distribution.
@@ 2 Preliminary results under SRD
The following lemma will be applied in the proof of the main result of this section, Theorem 2.2, deriving the asymptotic Gaussian distribution of in (21) under Specifically, Lemma 2.1 provides the asymptotic Gaussian distribution of the weighted periodogram operator under Its proof can be obtained in the same way as in Panaretos and Tavakoli (2013a), where this result is formulated for the separable Hilbert space
Lemma 2.1.
@@
Proof. See Theorem 3.7 in Panaretos and Tavakoli (2013a).
@@ 3 Second and fourth order bias asymptotics under LRD
This section provides new results on the bias asymptotics of the integrated empirical second and fourth order cumulant spectral density operators of under These results are applied in the derivation of Theorem 4.2 and Corollary 4.4, providing the consistency of the integrated weighted periodogram operator under In what follows, we assume and as
The rate of convergence to zero of the norm of the bias in the space of the of the integrated periodogram operator is obtained under LRD in the next lemma. The following well-known identity will be applied:
@@
The following corollary is obtained from Lemma 3.1, and provides the rate of convergence to zero of the bias of the integrated weighted periodogram operator, in the norm of the space under
Corollary 3.2.
@@
The rate of convergence to zero, in the norm of the space of the bias of the integrated empirical fourth–order cumulant spectral density operators of under LRD is obtained in Lemma 3.3 below. The following assumption is required:
@@ 4 A test for LRD in functional time series on
Consistency of the test based on is derived in this section. Specifically, Theorem 4.5 provides the almost surely divergence in the norm of the space of under The proof of this result follows from Proposition 4.1, showing the divergence in the norm of the space of the centering operator of and from Theorem 4.2 and Corollary 4.4, establishing the consistency of the integrated weighted periodogram operator under The implementation of the testing procedure in practice is also discussed.
Proposition 4.1.
@@
Theorem 4.2 implies the weak consistency of the integrated weighted periodogram operator under in the norm of the space
Corollary 4.4.
Under the conditions of Theorem 4.2, as
@@
Theorem 4.5.
Under assume that and that the bandwidth parameter for If conditions of Theorem 4.2 hold, then, as

### v4

@@ Abstract
M.D. Ruiz-Medina1 and Rosa M. Crujeiras2
1 University of Granada
2 University of Santiago de Compostela
MSC2020 subject classifications:
Primary. 60G10, 60G12, 60G18 , 60G20, 60G22
Secondary. 60G60
Keywords. Asymptotic normality; bias; compact manifolds; consistency; empirical cumulant spectral density operator; functional time series; integrated weighted periodogram operator; long-range dependence; spectral density operator.
@@ 1 Introduction
Spherical functional time series analysis helps in understanding the dynamics and spatiotemporal patterns of data that are embedded into the sphere, providing valuable insights for prediction, monitoring, and decision-making. Time series analysis of global temperature data distribution among other climate variables, usually arising in Climate Science and Meteorology, can be performed in a more efficient way by adopting a functional time series framework (see [38]). That is the case of ocean currents, and other marine functional time series to be analyzed in Oceanography studies (see, e.g., [41]; [43]). Other areas demanding this type of techniques are Geophysics, Astronomy and Astrophysics. In the last few decades, the cosmic microwave background radiation variation analysis over time has gained special attention (see [13]; [22]; [23]). In a more general manifold setting, functional time series analysis is often applied in Medical Imaging, Computer Vision and Graphics (see [42]; [45]; [44], among others). This paper focuses on the spectral analysis of functional time series in manifolds, with special attention to LRD analysis.
@@
The spectral analysis of functional time series has mainly been developed under Short Range Dependence (SRD). In this context, based on the weighted periodogram operator, a nonparametric framework is adopted in [28]. Particularly, the asymptotic normality of the functional discrete Fourier transform (fDFT), and the weighted periodogram operator of the curve data are proved under suitable summability conditions on the norm of the cumulant spectral density operators. The consistency of the weighted periodogram operator, in the Hilbert–Schmidt operator norm, is derived under SRD. In [29], a harmonic principal component analysis of functional time series in the temporal functional spectral domain is also obtained, based on a Karhunen–Loéve–like decomposition, the so–called Cramér–Karhunen–Loéve representation. In the context of functional regression, some applications are presented in [30], [34] and [39]. Hypothesis testing for detecting modelling differences in functional time series dynamics is addressed in [40] in the functional spectral domain.
@@
In LRD analysis of functional time series several problems still remain open. One of the key approaches in the current literature is presented in [19], where the eigendecomposition of the long-run covariance operator is considered, under an asymptotic semiparametric functional principal component framework. The consistent estimation of the dimension and the orthonormal functions spanning the dominant subspace, where the projected curve process displays the largest dependence range is derived. Fractionally integrated functional autoregressive moving averages processes constitute an interesting example of this modelling framework.
@@
A first attempt to characterize LRD in functional time series in the spectral domain can be found in [36], adopting the theoretical framework of operator–valued random fields, including fractional Brownian motion with operator–valued Hurst coefficient (see, e.g., [12]; [24]; [31] and [32]). The eigenvalues of the LRD operator are parameterized. These eigenvalues induce different levels of singularity at zero frequency, corresponding to different levels of temporal persistent of the process projected into different eigenspaces of the Laplace Beltrami operator. Under this LRD scenario, the integrated periodogram operator is proved to be asymptotically unbiased in the Hilbert–Schmidt operator norm. Minimum contrast estimation of the LRD operator is achieved in the spectral domain in a weak–consistent way under a Gaussian scenario. Interesting examples of this setting are analyzed in [26], where the spectral analysis of multifractionally integrated functional time series in manifolds is considered. In particular, multifractionally integrated spherical functional ARMA models (i.e., multifractionally integrated SPHARMA models) are analyzed through simulations. In this modelling framework, SRD and LRD can coexist at different spherical scales.
Up to our knowledge, no further developments have been achieved in the spectral analysis of LRD functional time series. Alternative contributions for stationary LRD functional sequences are based on the diagonalization of the heavy tail autocovariance kernel of the time–varying functional error term (see, e.g., [3]). However, under this modelling framework, functional spectral analysis can not be achieved in the time domain due to the assumed independence between the random components of the error term (see also [2]). Similar assertions hold for the two sample problem analyzed in [4]. In [20], a special family of LRD linear functional time series is analyzed with scalar coefficients displaying slow decay. Under stationarity, this LRD scenario constitutes a particular case of our framework when the elements of the spectral density operator family have degenerated pure point spectra corresponding to one infinite–dimensional eigenspace. The Cramér–Karhunen–Loéve representation above referred constitutes a powerful tool in the functional spectral analysis of weak-dependent functional time series (see [35]). In this paper, this representation is extended to our LRD stationary functional time series context, assuming the invariance of the cross-covariance kernels under the group of isometries of the manifold, given by a connected and compact two points homogeneous space.
@@
In this paper we perform a weighted periodogram operator based analysis, requiring the asymptotic analysis of the bias of the integrated empirical fourth–order cumulant spectral density operators, to prove consistency of the integrated weighted periodogram operator under LRD. Its application to spectral statistical hypothesis testing of LRD in –valued correlated sequences constitutes one of the main goals of this work. Here, denotes the space of real–valued square integrable functions on a Riemannian manifold embedded into given by a connected and compact two–point homogeneous space. The topological dimension of is and denotes the normalized Riemannian measure on In what follows, we will consider to be a functional sequence such that for every with denoting the probability measure defined on the basic probability space i.e., for every
@@
The invariance of the elements of the spectral density operator family of under the group of isometries of the manifold is assumed along the paper. A frequency–varying eigenvalue sequence then characterizes the pure point spectra of the elements of the spectral density operator family, with respect to the orthonormal basis of eigenfunctions of the Laplace–Beltrami operator. This invariance assumption is exploited in the derived Central Limit Theorem that characterizes the asymptotic distribution of the proposed test statistic operator under the null hypothesis, which states that displays SRD. In our formulation of the alternative hypothesis on LRD, we adopt a semiparametric framework in terms of a functional parameter given by the LRD operator. In contrast with the approach presented in [36], here we do not assume a parameterization of the eigenvalues of the LRD operator. Under this scenario, the rate of convergence to zero, in the corresponding norm, of the the bias of the integrated empirical second and fourth order cumulant spectral density operators is respectively obtained in Lemmas 2 and 3 under suitable conditions. Proposition 1 shows the divergence, in the Hilbert–Schmidt operator norm, of the mean of the test statistic operator under the alternative hypothesis. Theorem 2 derives suitable conditions, in particular, on the nonparametric functional spectral factor, to ensure consistency of the integrated weighted periodogram operator in the Hilbert–Schmidt operator norm under LRD. Theorem 3 then provides the almost surely divergence of the test statistic in the Hilbert–Schmidt operator norm under the alternative, yielding the consistency of the test.
@@
Theorem 3 also plays a crucial role in the implementation in practice of the proposed testing procedure, based on rejecting the null hypothesis when the random Fourier coefficients of the test operator statistic, suitably standardized according to Theorem 1, cross an upper or lower tail standard normal critical value. The orthonormal basis involved in the computation of these coefficients is constructed by the tensor product of the eigenfunctions of the Laplace Beltrami operator. The random projection methodology (see Theorem 4.1 in [8]) can be implemented here to alleviate the dimensionality problem. Specifically, when the moments of our test statistic operator satisfy the Carleman condition under the null hypothesis, our testing procedure is equivalent to rejecting the null hypothesis when the absolute value of a random projection of the test statistic is larger than an upper tail standard normal critical value. In the implementation of the random projection methodology, the Karhunen–Loéve expansion in Lemma 4 below can be considered for generation of the involved Gaussian random directions. In the simulation study undertaken, robust empirical sizes, and competitive values of the empirical power are displayed by our testing approach (see Section 5.4).
@@
The outline of the paper is as follows. In Section 1.1, the functional spectral background material is introduced. Our hypothesis testing procedure is formulated in a functional semiparametric spectral framework in Section 1.2. The asymptotic Gaussian distribution of the test statistic operator under the null hypothesis is obtained in Theorem 1 in Section 2. Asymptotics of the bias of the integrated empirical second and fourth order cumulant spectral density operators under LRD are derived in Section 3. Section 4 provides the preliminary results required for consistency of the test, which is derived in Theorem 3 of this section. Practical implementation is also discussed in Section 4. In Section 5.1, a simulation study is undertaken to illustrate the asymptotic Gaussian distribution of the proposed test statistic operator under the null hypothesis, in the context of SRD spherical functional time series. The consistency of the test is also illustrated in Section 5.2, in the framework of multifractionally integrated spherical functional time series. This numerical analysis is extended in Section 5.3 to a wider family of LRD operators allowing stronger persistency in time, displayed by the projected process in the dominant subspace. Section 5.4 analyzes empirical size and power properties of the test. Section 6 summarizes conclusions of the simulation study, focusing on the large functional sample size properties of our test statistic operator under different bandwidth parameter scenarios, from additional numerical results. The proofs of the results of this paper can be found in the Appendix.
@@ 1.1 Background
Along this work we will assume that in (1) is a stationary zero–mean functional sequence, with nuclear covariance operator family satisfying for every The elements of this family are characterized in the spectral domain by the spectral density operator family The assumed invariance of the elements of these families with respect to the group of isometries of yields to their diagonal series expansion in terms of with being the orthonormal basis of eigenfunctions of the Laplace–Beltrami operator on (see, e.g., [15]; [17]). In particular,
where, for every denotes the dimension of the eigenspace associated with the eigenvalue of the Laplace Beltrami operator (see, e.g., Section 2.1 in [21]). The equality means identity in the norm of the space of Hilbert–Schmidt operators on the space of complex–valued square integrable functions on Specifically, the equality in (2) means that
@@ 1.2 Hypothesis testing
Stationary SRD functional time series are characterized by the summability of the series of trace norms of the elements of the family of covariance operators (see, e.g., [28]). That is, displays SRD if and only if where denotes the space of trace operators on In our setting, this condition can be formulated as follows:
When (8) fails, is said to display LRD. In what follows, we will adopt the LRD scenario introduced in [36], given by
where the invariant positive self–adjoint operators and are composed yielding the definition of Specifically, denotes the LRD operator on Operator in (9) is interpreted as in [7], [31] and [32], where plays the role of operator–valued Hurst coefficient in the setting of fractional Brownian motion introduced in this framework. Moreover, is the regular spectral operator reflecting markovianess when the null space of coincides with Specifically, satisfies
where the operator integrals are understood as improper operator Stieltjes integrals which converge strongly (see, e.g., Section 8.2.1 in [33]).
We will apply the spectral theory of self–adjoint operators (see, e.g., [11]) in terms of the common spectral kernel
@@
where denotes the sequence of positive eigenvalues. For each is a continuous positive slowly varying function at in the Zygmund’s sense (see Definition 6.6 in [1], and Assumption IV in [36]). Equation (10) can be equivalently expressed, in terms of as
@@
for every (see [14] for the usual notion of convergence in the sense of generalized functions or distributions). Here, denotes the Dirac Delta distribution at zero frequency. Hence, in what follows we adopt the notation representing a truncated Dirac Delta distribution.
@@ 2 Preliminary results under SRD
The following lemma will be applied in the proof of the main result of this section, Theorem 1, deriving the asymptotic Gaussian distribution of in (21) under Specifically, Lemma 1 provides the asymptotic Gaussian distribution of the weighted periodogram operator under Its proof can be obtained in the same way as in [28], where this result is formulated for the separable Hilbert space
Lemma 1
@@
Proof. See Theorem 3.7 in [28].
@@ 3 Second and fourth order bias asymptotics under LRD
This section provides new results on the bias asymptotics in the Hilbert-Schmidt operator norm of the integrated empirical second and fourth order cumulant spectral density operators of under These results are applied in the derivation of Theorem 2 and Corollary 2, providing the consistency of the integrated weighted periodogram operator under In what follows, we assume and as
The rate of convergence to zero of the norm of the bias in the space of the integrated periodogram operator is obtained under LRD in the next lemma. The following well-known identity will be applied:
@@
The following corollary is obtained from Lemma 2, and provides the rate of convergence to zero of the bias of the integrated weighted periodogram operator, in the norm of the space under
Corollary 1
@@
The rate of convergence to zero, in the norm of the space of the bias of the integrated empirical fourth–order cumulant spectral density operators of under LRD is obtained in Lemma 3 below. The following assumption is required:
@@ 4 A test for LRD in functional time series on
Consistency of the test based on is derived in this section. Specifically, Theorem 3 provides the almost surely divergence in the norm of the space of under The proof of this result follows from Proposition 1, showing the divergence in the norm of the space of the centering operator of and from Theorem 2 and Corollary 2, establishing the consistency of the integrated weighted periodogram operator under The implementation of the testing procedure in practice is also discussed.
Proposition 1
@@
Theorem 2 implies the weak consistency of the integrated weighted periodogram operator under in the norm of the space
Corollary 2
Under the conditions of Theorem 2, as
@@
Theorem 3
Under assume that and that the bandwidth parameter for If conditions of Theorem 2 hold, then, as